Tensile strength tests were carried out on two different grades of wire rod, resulting in the accompanying data.

Grade Sample
Size
Sample
Mean
(kg/mm2)
Sample
SD
AISI 1064       m = 127     x = 106.9   s1 = 1.2
AISI 1078     n = 127 y = 126.8 s2 = 2.2

(a) Does the data provide compelling evidence for concluding that true average strength for the 1078 grade exceeds that for the 1064 grade by more than 10 kg/mm2? Test the appropriate hypotheses using a significance level of 0.01. State the relevant hypotheses.

H0: μ1064μ1078 = −10
Ha: μ1064μ1078 > −10H0: μ1064μ1078 = −10
Ha: μ1064μ1078 ≠ −10    H0: μ1064μ1078 = −10
Ha: μ1064μ1078 ≤ −10H0: μ1064μ1078 = −10
Ha: μ1064μ1078 ≥ −10H0: μ1064μ1078 = −10
Ha: μ1064μ1078 < −10


Calculate the test statistic and P-value. (Round your test statistic to two decimal places and your P-value to four decimal places.)

z =
P-value =


State the conclusion in the problem context.

Reject H0. The data does not suggest that the true average strength for the 1078 grade exceeds that for the 1064 grade by more than 10 kg/mm2.Fail to reject H0. The data suggests that the true average strength for the 1078 grade exceeds that for the 1064 grade by more than 10 kg/mm2.    Reject H0. The data suggests that the true average strength for the 1078 grade exceeds that for the 1064 grade by more than 10 kg/mm2.Fail to reject H0. The data does not suggest that the true average strength for the 1078 grade exceeds that for the 1064 grade by more than 10 kg/mm2.


(b) Estimate the difference between true average strengths for the two grades in a way that provides information about precision and reliability. (Use a 95% confidence interval. Round your answers to two decimal places.)

   ,  

kg/mm2
Answer #1

H0: μ1064μ1078 = −10
Ha: μ1064μ1078 < −10

pop 1 pop 2
sample mean x = 106.90 126.80
std deviation σ= 1.200 2.200
sample size n= 127 127
std error σx1-x2=√(σ21/n122/n2)    = 0.222
test stat z =(x1-x2o)/σx1-x2     = -44.52
p value : = 0.0000 (from excel:1*normsdist(-44.52)

  Reject H0. The data suggests that the true average strength for the 1078 grade exceeds that for the 1064 grade by more than 10 kg/mm2

b)

Point estimate of difference '=x1-x2 = -19.900
for 95 % CI value of z= 1.960 from excel:normsinv((1+0.95)/2)
margin of error E=z*std error = 0.436
lower bound=(x1-x2)-E = -20.34
Upper bound=(x1-x2)+E = -19.46
from above 95% confidence interval for population mean =(-20.34 , -19.46)